# Distributionally Robust Optimization with Adversarial Data Contamination

### v1

@@ Appendix D Auxiliary Results for Section 4
we define $L({\bm{w}},{\bm{\alpha}})=\frac{1}{N}\sum_{i=1}^{N}(\alpha^{i}{\bm{x}}_{i}\cdot{\bm{w}}-g_{i}(\alpha^{i}))+\psi({\bm{w}})$ and proceed to analyze Algorithm 1 as a primal-dual algorithm.
Upper bound for $a_{k}L({\bm{w}}_{k},{\bm{v}}_{\star})$.
@@
where the inequality follows from 1-strong convexity of $v\mapsto(v-\alpha_{k-1}^{i})^{2}$ and Fact B.1.
Lower bound for $a_{k}L({\bm{u}}_{\star},{\bm{\alpha}}_{k})$.
Let ${\bm{e}}_{k}={\mathcal{O}}({\bm{\beta}}_{k-1},\{{\bm{x}}_{i}\})-\frac{1}{N}\sum_{i=1}^{N}\beta_{k-1}^{i}{\bm{x}}_{i}$. It holds that
@@
where the last inequality follows from 1-strong convexity of ${\bm{u}}\mapsto\|{\bm{u}}-{\bm{w}}_{k-1}\|^{2}_{2}$, the construction ${\bm{w}}_{k}=\operatorname*{arg\,min}_{{\bm{u}}\in\mathbb{R}^{d}}\{\sum_{i=1}^{N}a_{k}({\bm{e}}_{k}+\frac{1}{N}\beta_{k-1}^{i}{\bm{x}}_{i})\cdot{\bm{u}}+a_{k}\psi({\bm{u}})+\frac{c_{k}}{2\gamma}\|{\bm{u}}-{\bm{w}}_{k-1}\|_{2}^{2}\}$, and Fact B.1. To continue the lower bound, note that $(c_{k})$ is chosen to be a decreasing sequence and we rewrite Equation 12 as follows:
@@
Plugging back in Equation 13 and using $c_{1}=2,c_{k}-c_{k+1}=1/T$, we have
@@
where the last inequality uses $\zeta$-boundedness of the domain of $g$ using Fact 2.4.
@@
Recall from Equation 11 that $\frac{1}{N}\sum_{i=1}^{N}\ell_{i}({\bm{x}}_{i}\cdot{\bm{w}})+\psi({\bm{w}})=\max_{{\bm{\alpha}}\in\mathbb{R}^{N}}L({\bm{w}},{\bm{\alpha}})$ for all ${\bm{w}}\in\mathbb{R}^{d}$. Setting ${\bm{v}}_{\star}=\operatorname*{arg\,max}_{{\bm{\alpha}}\in\mathbb{R}^{N}}L(\hat{\bm{w}}_{T},{\bm{\alpha}})$, we have $L(\hat{\bm{w}}_{T},{\bm{v}}_{\star})=\frac{1}{N}\sum_{i=1}^{N}\ell_{i}({\bm{x}}_{i}\cdot\hat{\bm{w}}_{T})+\psi(\hat{\bm{w}}_{T})$. Setting ${\bm{u}}_{\star}={\bm{w}}^{*}$, we have $L({\bm{u}}_{\star},\hat{\bm{\alpha}}_{T})\leq\frac{1}{N}\sum_{i=1}^{N}\ell_{i}({\bm{x}}_{i}\cdot{\bm{w}}^{*})+\psi({\bm{w}}^{*})$. Hence,
@@
Corollary D.1.
Under the same setup as in Proposition 4.1, for every iteration $k\leq T$, it holds that
$\|{\bm{w}}_{k}-{\bm{w}}^{*}\|_{2}\leq 4\|{\bm{w}}_{0}-{\bm{w}}^{*}\|_{2}.$
Therefore, we have $\|{\bm{w}}_{k}\|_{2}-\|{\bm{w}}^{*}\|_{2}\leq 4\|{\bm{w}}_{0}\|_{2}+4\|{\bm{w}}^{*}\|_{2}$ and thus
$\|{\bm{w}}_{k}\|_{2}\leq 4\|{\bm{w}}_{0}\|_{2}+5\|{\bm{w}}^{*}\|_{2}.$
In particular, the norm of the convex combination also satisfies
$\|\hat{\bm{w}}_{T}\|_{2}=\Big{\|}\frac{1}{A_{T}}\sum_{k=1}^{N}a_{k}{\bm{w}}_{k}\Big{\|}_{2}=O(\|{\bm{w}}_{0}\|_{2}+\|{\bm{w}}^{*}\|_{2}).$ (17)
@@
Equation 14 holds for all $k\in\mathbb{Z}_{+}$, in particular for all $k\leq T=2\zeta\sqrt{\|{\bm{\Sigma}}_{B}\|_{\operatorname{op}}}/\Delta$. Fix $k\leq T$.
Denote $D_{0}=\|{\bm{u}}_{\star}-{\bm{w}}_{0}\|_{2}$. Plugging $\gamma=D_{0}/(\zeta\sqrt{N})$ into Equation 14 gives
@@
Setting ${\bm{u}}_{\star}={\bm{w}}^{*}$ and ${\bm{v}}_{\star}=\operatorname*{arg\,max}_{{\bm{\alpha}}\in\mathbb{R}^{N}}L(\hat{\bm{w}}_{k},{\bm{\alpha}})$, it holds that $L(\hat{\bm{w}}_{k},{\bm{v}}_{\star})-L({\bm{u}}_{\star},\hat{\bm{\alpha}}_{k})\geq 0$, because it is the difference of the function value of the problem (P) at $\hat{\bm{w}}_{k}$ and the optimal solution ${\bm{w}}^{*}$. Thus,
$3\zeta D_{0}\sqrt{N}+\frac{D_{0}{k}^{2}\sqrt{N}\Delta^{2}}{2\zeta\|{\bm{\Sigma}}_{\bm{B}}\|_{\operatorname{op}}}-\frac{\zeta\sqrt{N}}{2D_{0}}\|{\bm{w}}_{k}-{\bm{u}}_{\star}\|_{2}^{2}\geq 0,$
which simplifies to
@@
D.2 Proof of Corollary 4.3
We restate Corollary 4.3 for completeness: See 4.3
@@
Let $\bar{F}({\bm{w}})=\mathbb{E}_{({\bm{x}},y)\sim{\mathbb{P}}_{0}}[\ell({\bm{w}};{\bm{x}},y)]+\rho\zeta\|{\bm{w}}\|_{s}$. By Equation 3, this is the population $W_{1}$-DRO objective: $\max_{{\mathbb{Q}}:W_{1}^{c}({\mathbb{P}}_{0},{\mathbb{Q}})\leq\rho}\mathbb{E}_{{\bm{\xi}}\sim{\mathbb{Q}}}[\ell({\bm{w}};{\bm{\xi}})]$. Let ${\bar{\bm{w}}}^{*}=\operatorname*{arg\,min}_{{\bm{w}}:\|{\bm{w}}\|_{2}\leq W}\bar{F}({\bm{w}})$.
Let $F^{\prime}_{S}({\bm{w}})=\frac{1}{|S^{\prime}|}\sum_{i:\tilde{\bm{x}}_{i}\in S^{\prime}}\ell_{y_{i}}({\bm{x}}_{i}\cdot{\bm{w}})+\rho\zeta\|{\bm{w}}\|_{s}$ be the empirical objective based on the set of stable covariates, denoted by $S^{\prime}$, whose existence is guaranteed with high probability by Fact 2.8. To avoid confusion, we highlight that ${\bm{w}}^{*}$ is defined to be the minimizer of $F^{\prime}_{S}({\bm{w}})$.
Theorem 4.2 implies that Algorithm 1, when run with ${\bm{w}}_{0}={\bm{0}}$ on $\epsilon$-corrupted data, outputs ${\bm{w}}_{T}$ such that with probability at least $0.995$:
Using Fact B.2, with probability at least $0.995$:
$\sup_{{\bm{w}}:\|{\bm{w}}\|_{2}\leq W}|\bar{F}({\bm{w}})-\hat{F}_{N}({\bm{w}})|=O(\zeta W\sqrt{d/N}).$ (19)
Combining these results with a union bound, with probability at least $0.99$:
$\bar{F}({\bm{w}}_{T})$ (19)${{\leq}}\hat{F}_{N}({\bm{w}}_{T})+O(\zeta W\sqrt{d/N})$ )
@@
(18)${{\leq}}\left(\min_{{\bm{w}}^{\prime}:\|{\bm{w}}^{\prime}\|_{2}\leq W}\hat{F}_{N}({\bm{w}}^{\prime})\right)+O(\sigma\zeta\sqrt{\epsilon}\|{\bm{w}}^{*}\|_{2})+O(\zeta W\sqrt{d/N})$ )
$\leq\hat{F}_{N}({\bar{\bm{w}}}^{*})+O(\sigma\zeta\sqrt{\epsilon}\|{\bm{w}}^{*}\|_{2})+O(\zeta W\sqrt{d/N})$
(19)${{\leq}}\bar{F}({\bar{\bm{w}}}^{*})+O(\sigma\zeta\sqrt{\epsilon}\|{\bm{w}}^{*}\|_{2})+O(\zeta W\sqrt{d/N})\;.$ ) .
The term $\bar{F}({\bar{\bm{w}}}^{*})$ is $\inf_{{\bm{w}}:\|{\bm{w}}\|_{2}\leq W}\bar{F}({\bm{w}})$. The excess risk is bounded by $C_{A}\sigma\zeta\sqrt{\epsilon}W+2O(\zeta W\sqrt{d/N})$. Plugging the sample size $N=O(\zeta W^{2}d\log(d)/\epsilon)$ into the above error bound, the uniform convergence error term $O(\zeta W\sqrt{d/N})$ becomes:
$O(\zeta W\sqrt{d/N})=O\left(\zeta W\sqrt{\frac{d}{W^{2}d\log d/\epsilon}}\right)=O(\zeta\sqrt{\epsilon}/\log d)=O(\zeta\sqrt{\epsilon}),$
D.3 Further Details for SVM Example

### v2

@@ 4 $A_{k}=a_{k}+A_{k-1},c_{k}=c_{k-1}-1/T$, where $a_{k}=\sqrt{N}/\sigma$
we define $L({\bm{w}},{\bm{\alpha}})=\frac{1}{N}\sum_{i=1}^{N}(\alpha^{i}{\bm{x}}_{i}\cdot{\bm{w}}-g_{i}(\alpha^{i}))+\psi({\bm{w}})$ and proceed to analyze Algorithm˜1 as a primal-dual algorithm.
First, we show an upper bound for $a_{k}L({\bm{w}}_{k},{\bm{v}}_{\star})$ for any given $\bm{v}$$_{\star}$ to be chosen later.
@@
where the inequality follows from 1-strong convexity of $v\mapsto(1/2)(v-\alpha_{k-1}^{i})^{2}$ and Fact˜4.2.
Next, we prove the lower bound for $a_{k}L({\bm{u}}_{\star},{\bm{\alpha}}_{k})$. Let ${\bm{e}}_{k}={\mathcal{O}}({\bm{\beta}}_{k-1},\{{\bm{x}}_{i}\})-\frac{1}{N}\sum_{i=1}^{N}\beta_{k-1}^{i}{\bm{x}}_{i}$. It holds that
@@
where the last inequality follows from 1-strong convexity of ${\bm{u}}\mapsto\frac{1}{2}\|{\bm{u}}-{\bm{w}}_{k-1}\|^{2}_{2}$, the construction ${\bm{w}}_{k}=\operatorname*{arg\,min}_{{\bm{u}}\in\mathbb{R}^{d}}\{\sum_{i=1}^{N}a_{k}({\bm{e}}_{k}+\frac{1}{N}\beta_{k-1}^{i}{\bm{x}}_{i})\cdot{\bm{u}}+a_{k}\psi({\bm{u}})+\frac{c_{k}}{2\gamma}\|{\bm{u}}-{\bm{w}}_{k-1}\|_{2}^{2}\}$, and Fact˜4.2. To continue the lower bound, note that $(c_{k})$ is chosen to be a decreasing sequence and we rewrite Equation˜9 as follows:
@@
Plugging back in Equation˜10 and using $c_{1}=2,c_{k}-c_{k+1}=1/T$, we have
@@
where the last inequality uses $\zeta$-boundedness of the domain of $g$ using Fact˜2.4.
@@
Recall from Equation˜8 that $\frac{1}{N}\sum_{i=1}^{N}\ell_{i}({\bm{x}}_{i}\cdot{\bm{w}})+\psi({\bm{w}})=\max_{{\bm{\alpha}}\in\mathbb{R}^{N}}L({\bm{w}},{\bm{\alpha}})$ for all ${\bm{w}}\in\mathbb{R}^{d}$. Setting ${\bm{v}}_{\star}=\operatorname*{arg\,max}_{{\bm{\alpha}}\in\mathbb{R}^{N}}L(\hat{\bm{w}}_{T},{\bm{\alpha}})$, we have $L(\hat{\bm{w}}_{T},{\bm{v}}_{\star})=\frac{1}{N}\sum_{i=1}^{N}\ell_{i}({\bm{x}}_{i}\cdot\hat{\bm{w}}_{T})+\psi(\hat{\bm{w}}_{T})$. Setting ${\bm{u}}_{\star}={\bm{w}}^{*}$, we have $L({\bm{u}}_{\star},\hat{\bm{\alpha}}_{T})\leq\frac{1}{N}\sum_{i=1}^{N}\ell_{i}({\bm{x}}_{i}\cdot{\bm{w}}^{*})+\psi({\bm{w}}^{*})$. Hence,
@@
Note that Equation˜11 directly implies boundedness of iterates:
Corollary 4.3.
Under the same setup as in Proposition˜4.1, for every iteration $k\leq T$, it holds that $\|{\bm{w}}_{k}-{\bm{w}}^{*}\|_{2}\leq 4\|{\bm{w}}_{0}-{\bm{w}}^{*}\|_{2}.$ Therefore, we have $\|{\bm{w}}_{k}\|_{2}-\|{\bm{w}}^{*}\|_{2}\leq 4\|{\bm{w}}_{0}\|_{2}+4\|{\bm{w}}^{*}\|_{2}$ and thus $\|{\bm{w}}_{k}\|_{2}\leq 4\|{\bm{w}}_{0}\|_{2}+5\|{\bm{w}}^{*}\|_{2}.$ In particular, the norm of the convex combination also satisfies
$\|\hat{\bm{w}}_{k}\|_{2}=\Big\|\frac{1}{A_{T}}\sum_{k=1}^{N}a_{k}{\bm{w}}_{k}\Big\|_{2}=O(\|{\bm{w}}_{0}\|_{2}+\|{\bm{w}}^{*}\|_{2}).$ (14)
@@
Equation˜11 holds for all $k\in\mathbb{Z}_{+}$, in particular for all $k\leq T=2\zeta\sqrt{\|{\bm{\Sigma}}_{B}\|_{\operatorname{op}}}/\Delta$. Fix $k\leq T$.
Denote $D_{0}=\|{\bm{u}}_{\star}-{\bm{w}}_{0}\|_{2}$. Plugging $\gamma=D_{0}/(\zeta\sqrt{N})$ into Equation˜11 gives
@@
Setting ${\bm{u}}_{\star}={\bm{w}}^{*}$ and ${\bm{v}}_{\star}=\operatorname*{arg\,max}_{{\bm{\alpha}}\in\mathbb{R}^{N}}L(\hat{\bm{w}}_{k},{\bm{\alpha}})$, it holds that $L(\hat{\bm{w}}_{k},{\bm{v}}_{\star})-L({\bm{u}}_{\star},\hat{\bm{\alpha}}_{k})\geq 0$, because it is the difference of the function value of the problem (P) at $\hat{\bm{w}}_{k}$ and the optimal solution $\bm{w}$$^{*}$. Thus, $3\zeta D_{0}\sqrt{N}+\frac{D_{0}{k}^{2}\sqrt{N}\Delta^{2}}{2\zeta\|{\bm{\Sigma}}_{\bm{B}}\|_{\operatorname{op}}}-\frac{\zeta\sqrt{N}}{2D_{0}}\|{\bm{w}}_{k}-{\bm{u}}_{\star}\|_{2}^{2}\geq 0,$ which simplifies to
@@
We highlight that the error in Proposition˜4.1 has dependence on the initial distance $\|{\bm{w}}_{0}-{\bm{w}}^{*}\|_{2}$ rather than any notion of weight domain radius. This error is optimal up to a constant; see Lemma˜4.8.
Theorem 4.4.
Suppose the loss function satisfies Assumption˜1.3. A number of $N$ samples are generated per Assumption˜3.8 and then $\epsilon$-corrupted per Definition˜1.1. With probability at least $1-\tau$, Algorithm˜1 outputs $\hat{\bm{w}}_{T}$ solving (1) with error $O(\sigma\zeta\sqrt{\epsilon}(\|{\bm{w}}^{*}\|_{2}+\|{\bm{w}}_{0}\|_{2}))$ in the objective function for some sample size $N=O((d\log d+\log(1/\tau))/\epsilon)$.
@@
By Assumption˜3.8 and Lemma˜B.2, the uncorrupted version of covariates $S=\{{\bm{x}}_{i}\}_{i=1}^{N}$ is generated from a distribution of covariance $\Sigma$ with $\|\Sigma\|_{\operatorname{op}}\leq\sigma^{2}$. By Fact˜2.8, with probability at least $1-\tau$, the set $S^{\prime}=\{{\bm{x}}\in S:\|{\bm{x}}-{\bm{\mu}}_{S}\|_{2}\leq 2\sigma\sqrt{d/\epsilon}\}$ satisfies $(\epsilon,O(\sqrt{\epsilon}))$-stability and $|S^{\prime}|\geq(1-\epsilon)|S|$. Let $T$ be the given $\epsilon$-corrupted version of $S$. Because $|S^{\prime}|\geq 1-\epsilon$, we can view $T$ as a $2\epsilon$-corrupted version of $S^{\prime}$.
The key idea is that although the algorithm is run on the corrupted samples, its analysis tracks quantities related to the stable samples. More precisely, we establish a coupling argument that all iterates in Algorithm˜1 (on input corrupted samples $T$ and corresponding labels) are the same as the corresponding iterates in Algorithm˜2 (with input samples replaced by stable covariates $S^{\prime}$ and corresponding labels), where ${\bm{z}}_{k}$ in Algorithm˜2 is chosen to be $\bar{\bm{z}}_{k}$. For $\bar{\bm{z}}_{k}$ to be a valid choice, we will verify that $\|\bar{\bm{z}}_{k}-\frac{1}{N}\sum_{i=1}^{N}\beta_{k-1}^{i}{\bm{x}}_{i}^{s}\|_{2}\leq\Delta$ for all $k$.
Let ${\mathcal{I}}_{\text{bad}}$ be the indices of the samples in $T\setminus S^{\prime}$, which include corrupted samples and samples with large covariates. We proceed with a proof by induction. For iteration $k=1$, both algorithms are initialized at the same point ${\bm{w}}_{0}$. Since ${\bm{\beta}}_{0}=\bar{\bm{\beta}}_{0}$, Proposition˜3.10 verifies that $\|\bar{\bm{z}}_{k}-\frac{1}{N}\sum_{i=1}^{N}\beta_{k-1}^{i}{\bm{x}}_{i}^{s}\|_{2}\leq\Delta$ for $k=1$. Because only samples of indices in ${\mathcal{I}}_{\text{bad}}$ are different in the input of both algorithms, $\bar{\bm{\alpha}}_{1}$ is equal to ${\bm{\alpha}}_{1}$ except for indices in ${\mathcal{I}}_{\text{bad}}$.
We inductively assume that both algorithms produce the same iterates at iteration k1k-1, and $\bar{\bm{\alpha}}_{k-1}$ is equal to ${\bm{\alpha}}_{k-1}$ except for indices in ${\mathcal{I}}_{\text{bad}}$. The key observation here is that quantities $\bar{\bm{\alpha}}$, $\bar{\bm{\beta}}$, and ${\bm{x}}^{c}$ at the start of iteration $k$ of Algorithms˜1 and 2 only differ in indices ${\mathcal{I}}_{\text{bad}}$. Moreover, both $\beta_{k}^{i}$ and $\bar{\beta}_{k}^{i}$ are $3\zeta$-uniformly bounded for all $i$ and $k$. Therefore, both $\{\beta^{i}_{k-1}{\bm{x}}^{c}_{i}\}_{i\in[N]}$ and $\{\bar{\beta}^{i}_{k-1}{\bm{x}}^{c}_{i}\}_{i\in[N]}$ are $2\epsilon$-corrupted versions of the stable set $\{\beta^{i}_{k-1}{\bm{x}}^{s}_{i}\}$ with the same stability parameters. Then Proposition˜3.10 verifies again that $\|\bar{\bm{z}}_{k}-\frac{1}{N}\sum_{i=1}^{N}\beta_{k-1}^{i}{\bm{x}}_{i}^{s}\|_{2}\leq\Delta$ for iteration $k$, which implies that both algorithms produce the same iterate ${\bm{w}}_{k}$, completing the induction.
It remains to analyze the convergence guarantee of Algorithm˜2. Fact˜2.6 and Fact˜2.8 immediately imply that $\|{\bm{\Sigma}}_{\bm{B}}\|_{\operatorname{op}}=O(\sigma^{2})$, where ${\bm{\Sigma}}_{\bm{B}}$ is the empirical covariance matrix of $S^{\prime}$. Applying Proposition˜4.1 with $\Delta=O(\sigma\zeta\sqrt{\epsilon})$ implies that for stable set $\{{\bm{x}}_{i}^{s}\}_{i=1}^{N}$, $\frac{1}{N}\sum_{i=1}^{N}l_{i}({\bm{x}}_{i}^{s}\cdot\hat{\bm{w}}_{T})+\psi(\hat{\bm{w}}_{T})\leq\frac{1}{N}\sum_{i=1}^{N}l_{i}({\bm{x}}_{i}^{s}\cdot{\bm{w}}^{*})+\psi({\bm{w}}^{*})+3\|{\bm{w}}^{*}-{\bm{w}}_{0}\|_{2}\Delta$. We conclude the proof by noting that for all ${\bm{w}}\in\mathbb{R}^{d}$, $|\frac{1}{N}\sum_{i=1}^{N}l_{i}({\bm{x}}_{i}^{s}\cdot{\bm{w}})-\frac{1}{N}\sum_{i=1}^{N}l_{i}({\bm{x}}_{i}\cdot{\bm{w}})|=O(\zeta\sigma\|{\bm{w}}\|_{2}\sqrt{\epsilon})$. ∎
By standard uniform convergence results and initializing at ${\bm{w}}_{0}={\bm{0}}$, we get the following corollary whose proof can be found in Appendix˜C.
Corollary 4.5.
Suppose the loss function satisfies Assumption˜1.3 and $N$ samples are generated according to Assumption˜3.8 and then $\epsilon$-corrupted. Let $\tau\in(0,1)$ and $c(({\bm{x}},y),({\bm{x}}^{\prime},y^{\prime}))=\|{\bm{x}}-{\bm{x}}^{\prime}\|_{r}+\chi(y-y^{\prime})$. If in addition we assume that $\|{\bm{x}}\|_{2}\leq C\sigma\sqrt{d}$ for some constant $C>0$, then there exists a sample size $N=O(W^{2}d(\log(d)+\log(1/\tau))/\epsilon)$ such that with probability at least $1-\tau$, Algorithm˜1 outputs ${\bm{w}}_{T}$ that satisfies
$\max_{{\mathbb{Q}}:W_{1}^{c}({\mathbb{P}}_{0},{\mathbb{Q}})\leq\rho}\mathbb{E}_{{\mathbb{Q}}}[\ell({\bm{w}}_{T};{\bm{\xi}})]\leq\inf_{{\bm{w}}:\|{\bm{w}}\|_{2}\leq W}\max_{{\mathbb{Q}}:W_{1}^{c}({\mathbb{P}}_{0},{\mathbb{Q}})\leq\rho}\mathbb{E}_{{\mathbb{Q}}}[\ell({\bm{w}};{\bm{\xi}})]+O(\sigma\zeta\sqrt{\epsilon}\|{\bm{w}}^{*}\|_{2}).$
Remark 4.6.
We remark that our algorithmic framework can be adapted to solve the empirical version of (7) in Example˜3.7, which takes the form $\min_{\bm{w}}\left(\sqrt{\frac{1}{N}\sum_{i=1}^{N}({\bm{x}}_{i}\cdot{\bm{w}}-y_{i})^{2}}+\psi({\bm{w}})\right)$, where $\psi({\bm{w}})=\rho\|{\bm{w}}\|_{s}$. Instead of using Fenchel conjugacy on separable loss terms, we can employ the self-duality of $\ell_{2}$ norm. Let ${\bm{v}}({\bm{w}})\in\mathbb{R}^{N}$ be the vector of residuals with entries $v_{i}({\bm{w}})={\bm{x}}_{i}\cdot{\bm{w}}-y_{i}$. The loss term can be written as $\frac{1}{\sqrt{N}}\|{\bm{v}}({\bm{w}})\|_{2}=\frac{1}{\sqrt{N}}\sup_{{\bm{z}}\in\mathbb{R}^{N}:\|{\bm{z}}\|_{2}\leq 1}{\bm{v}}({\bm{w}})^{\top}{\bm{z}}$. This yields the min-max problem $\min_{{\bm{w}}}\max_{{\bm{z}}\in\mathbb{R}^{N}:\|{\bm{z}}\|_{2}\leq 1}\left\{\frac{1}{\sqrt{N}}\sum_{i=1}^{N}z_{i}({\bm{x}}_{i}\cdot{\bm{w}}-y_{i})+\psi({\bm{w}})\right\}$. This structure is amenable to Algorithm˜1 with two modifications. First, the primal update (Line 1) requires a robust estimate of the vector $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\bar{z}_{k-1}^{i}{\bm{x}}_{i}$. Since the constraint $\|{\bm{z}}\|_{2}\leq 1$ implies that each weight $z_{i}$ is bounded by $|z_{i}|\leq 1$, this weighted sum can be computed by our robust oracle per Proposition˜3.10. Second, the dual update (Line 1) is no longer separable but becomes a joint proximal update for the vector $\bm{z}$}, which corresponds to an efficient projection onto the Euclidean unit ball.
Finally, we argue that the error in Theorem˜4.4 is optimal (up to constant factors). We rely on the following lower bound of robust mean estimation for distributions with bounded variance, which can be found in e.g., [DK23a, Lemma 1.11 ].
Fact 4.7 (Lower Bound for Robust Mean Estimation).
Let $\mathbb{D}$} be the family of all one-dimensional distributions with variance at most $\sigma^{2}$. Let $\epsilon\in(0,1/2)$. Any algorithm with access to $\epsilon$-corrupted samples from an unknown distribution ${\mathbb{P}}\in{\mathbb{D}}$ incurs an additive error $\Omega(\sigma\sqrt{\epsilon})$ in the worst case to estimate $\mathbb{E}[{\mathbb{P}}]$.
The lower bound is established by constructing two “hard” distributions, ${\mathbb{P}}_{A}$ and ${\mathbb{P}}_{B}$, which an adversary can make indistinguishable. Specifically, let ${\mathbb{P}}_{A}(0)=1-\epsilon$ and ${\mathbb{P}}_{A}(\sigma/\sqrt{\epsilon})=\epsilon$ and ${\mathbb{P}}_{B}(0)=1-\epsilon$ and ${\mathbb{P}}_{B}(-\sigma/\sqrt{\epsilon})=\epsilon$. It is easy to verify that $\operatorname{Var}({\mathbb{P}}_{A})=\operatorname{Var}({\mathbb{P}}_{B})=\sigma^{2}(1-\epsilon)<\sigma^{2}$ and
$d_{\mathrm{TV}}({\mathbb{P}}_{A},{\mathbb{P}}_{B})$ $=\frac{1}{2}\sum_{x\in\{-\sigma/\sqrt{\epsilon},\;0,\;\sigma/\sqrt{\epsilon}\}}|{\mathbb{P}}_{A}(x)-{\mathbb{P}}_{B}(x)|$
$=\frac{1}{2}(|0-\epsilon|+|(1-\epsilon)-(1-\epsilon)|+|\epsilon-0|)=\epsilon.$
By [DK23a, Proposition 1.7 ], if $d_{\mathrm{TV}}({\mathbb{P}}_{A},{\mathbb{P}}_{B})\leq 2\epsilon$, an adversary can make them indistinguishable under $\epsilon$-contamination by outputting contaminated distribution $({\mathbb{P}}_{A}+{\mathbb{P}}_{B})/2$.
Now we establish the lower bound for the optimality gap based on Fact˜4.7.
Lemma 4.8 (Lower Bound on Optimality Gap).
Let $\epsilon\in(0,1/2)$ and $D>0$. For any $\zeta>0$ and $\sigma>0$, there exists a $\zeta$-Lipschitz generalized linear loss $\ell({\bm{w}};{\bm{x}})$ and two one-dimensional distributions, ${\mathbb{P}}_{A}$ and ${\mathbb{P}}_{B}$, both with variance at most $\sigma^{2}$ (as in Fact˜4.7), such that any algorithm given $\epsilon$-corrupted samples from an unknown distribution ${\mathbb{P}}\in\{{\mathbb{P}}_{A},{\mathbb{P}}_{B}\}$ and returning $\hat{{\bm{w}}}$ for the problem $\min_{{\bm{w}}\in[-D/2,D/2]}\mathbb{E}_{{\mathbb{P}}}[\ell({\bm{w}};{\bm{x}})]$ must incur an optimality gap
@@
$\max_{{\mathbb{P}}\in\{{\mathbb{P}}_{A},{\mathbb{P}}_{B}\}}\left\{\mathbb{E}_{{\mathbb{P}}}[\ell(\hat{{\bm{w}}};{\bm{x}})]-\min_{{\bm{w}}\in[-D/2,D/2]}\mathbb{E}_{{\mathbb{P}}}[\ell({\bm{w}};{\bm{x}})]\right\}=\Omega\big(\zeta D\sigma\sqrt{\epsilon}\big).$
Proof.
The proof is a reduction from the one-dimensional robust mean estimation lower bound. We adopt the hard instance distributions ${\mathbb{P}}_{A}$ and ${\mathbb{P}}_{B}$ from Fact˜4.7.
Consider the one-dimensional data space $\bm{x}$$\in\mathbb{R}$ and the parameter interval ${\bm{w}}\in[-D/2,D/2]$, which has diameter $D$. We choose the linear loss function $\ell({\bm{w}};{\bm{x}})=-\zeta{\bm{w}}\cdot{\bm{x}}$. This loss is convex in $\bm{w}$} and $\zeta$-Lipschitz in ${\bm{w}}\cdot{\bm{x}}$. The population objective for a distribution $\mathbb{P}$} with mean $\mu$ is
$\mathbb{E}_{{\bm{x}}\sim{\mathbb{P}}}[\ell({\bm{w}};{\bm{x}})]=-\zeta{\bm{w}}\mathbb{E}_{{\mathbb{P}}}[{\bm{x}}]=-\zeta\mu{\bm{w}}.$
The unique minimizer of this objective over ${\bm{w}}\in[-D/2,D/2]$ is ${\bm{w}}^{*}=(D/2)\cdot\operatorname{sign}(\mu)$.
We now consider the two hard distributions ${\mathbb{P}}_{A}$ and ${\mathbb{P}}_{B}$ from Fact˜4.7, with means $\mu_{A}=+\sigma\sqrt{\epsilon}$ and $\mu_{B}=-\sigma\sqrt{\epsilon}$ and the corresponding optimum ${\bm{w}}^{*}_{A}=+D/2$ and ${\bm{w}}^{*}_{B}=-D/2$.
By Fact˜4.7, an adversary can provide an $\epsilon$-corrupted dataset to the algorithm that is statistically indistinguishable from a corrupted sample of ${\mathbb{P}}_{A}$ or ${\mathbb{P}}_{B}$. Let the algorithm’s output on this corrupted dataset be $\hat{{\bm{w}}}\in[-D/2,D/2]$.
We now show that $\hat{{\bm{w}}}$ has a large optimality gap for at least one of these two distributions. We compute the gap for both cases:
[…]
Similarly, we have
[…]
$=\zeta(-\mu_{B})(\hat{{\bm{w}}}+D/2)=\zeta\mu_{A}(\hat{{\bm{w}}}+D/2)\qquad(\text{since }-\mu_{B}=\mu_{A}).$
[…]
The minimum of $)\max(\mathrm{Gap}_{A}(\hat{{\bm{w}}}),\mathrm{Gap}_{B}(\hat{{\bm{w}}}$)) is thus $\mathrm{Gap}_{A}(0)=\zeta\mu_{A}(D/2-0)=\zeta(c\sigma\sqrt{\epsilon})(D/2)=\Omega(\zeta D\sigma\sqrt{\epsilon}).$ Since any output $\hat{{\bm{w}}}$ must have a gap of at least this much in one of the two indistinguishable cases, the lower bound holds for any algorithm. ∎
4.1 Application to SVMs: Removing Dependence on the Optimal Solution Norm
The error bound in Corollary˜4.5 depends on the norm of the optimal solution, $\|{\bm{w}}^{*}\|_{2}$, which may be large or unknown. We now show that this dependence can be removed for certain structured problems by leveraging stronger distributional assumptions beyond a bounded covariance.
We present a key application to Support Vector Machine (SVM) classification with the hinge loss. By assuming the covariate distribution satisfies a standard anti-concentration property, we can provably restrict the search space for the model weights. This allows us to establish an error rate $O(\epsilon^{1/4})$, independent of $\|{\bm{w}}^{*}\|_{2}$. This subsection, following a similar line of arguments to [Dia+19a, Section 3 ], provides the necessary assumptions and detailed analysis for this result.